OpenStudy (anonymous):

y / y^2-y-20 - 2 / y+4 HELP APPRECIATED

OpenStudy (btaylor):

factor the first denominator (y^2 - y - 20) into (y-5)(y+4) now, do you see what you need to do to get a common denominator?

OpenStudy (anonymous):

how do you get rid of the y^2

OpenStudy (saifoo.khan):

PLease use brackets. @j-free

OpenStudy (anonymous):

is this your expression? \[\huge \frac{y}{y^2-y-20}-\frac{2}{y+4} \]

OpenStudy (saifoo.khan):

that's HUGE.

OpenStudy (anonymous):

(y/y^2-y-20) - (2/y+4)

OpenStudy (anonymous):

yes... it is... among other things... :)

OpenStudy (anonymous):

@dpaInc YES YES

OpenStudy (saifoo.khan):

Indeed. ;D

OpenStudy (anonymous):

@j-free , started you off with factorizing the denominator of that first rational expression. can you get that second fraction to have the common denominator?

OpenStudy (anonymous):

i meant @BTaylor ... ^^^

OpenStudy (anonymous):

2(y+2)

OpenStudy (btaylor):

not quite. you need to multiply the entire second term by (y-5)/(y-5)

OpenStudy (anonymous):

y+4/y-5

OpenStudy (anonymous):

y+20/y+25

OpenStudy (btaylor):

\[\frac{y}{(y+4)(y-5)}-\frac{2(y-5)}{(y+4)(y-5)}\] now combine the numerators...

OpenStudy (anonymous):

y-5 right?

OpenStudy (btaylor):

no, you get as your numerator y-2y+10 --> 10-y I believe that is as simple as you can make it.

OpenStudy (anonymous):

dont the denominators just cancel out now?

OpenStudy (btaylor):

no, the final answer is \[\frac{10-y}{(y+4)(y-5)}\] I just didn't include the denominator in my last response.

OpenStudy (anonymous):

10-y / y^2-20

OpenStudy (btaylor):

no, 10-y / y^2-y-20